Maria is now 3 years younger than her sister, Ann. Four years ago, Maria's age was one-half Ann's age. How old s each girl now?
Answer by DrBeeee(684)

ALWAYS define variables to represent the unknowns.
THEN write equation(s) relating the variables to the problem statements.
THEN solve and check your answer.
Plesae do these steps to LEARN how to solve problems using ALGEBRA.
Ley A = Ann's present age
Let M = Maria's present age
The first sentence says
(1) M = A - 3
The second statemnt says
(2) M - 4 = (A - 4)/2
Do you agree?
Now solve (1) and (2) using substitution. Place (1) into (2) to get
(3) (A - 3) - 4 = (A - 4)/2 or
(4) A - 7 = (A - 4)/2
Now multiply both sides by 2 to get
(5) 2A - 14 = A - 4 or
(6) A = 14 - 4 or
(7) A = 10
Using (1) we get
(8) M = 10 - 3 or
(9) M = 7
Check the answers using (2).
Is (7 - 4 = (10 - 4)/2)?
Is (3 = 6/2)?
Is (3 = 3)? Yes
Answer: At the present time Ann is 10 years old and Maria is 7 years old.